# Swiss Army Knife: Synergizing BiAses in Knowledge from Vision Foundation Models for Multi-Task Learning

### v1

@@ D.2 Vision Foundation Models
Vision Foundation Models (VFMs) are large-scale, general-purpose vision models trained on massive datasets, demonstrating exceptional performance across a wide range of downstream tasks, particularly when generalizing to unseen tasks and domains. Some VFMs are trained with specific tasks, such as ViT (Dosovitskiy et al., 2021), DeiT (Touvron et al., 2021), and Swin Transformer (Liu et al., 2021c) for image classification, Segment Anything Model (SAM) (Kirillov et al., 2023; Ravi et al., 2024) for promptable segmentation, and Depth Anything (Yang et al., 2024b; c) for monocular depth estimation. Others are trained with pretext tasks agnostic to downstream tasks. For instance, CLIP (Radford et al., 2021) and its derivatives (Zhai et al., 2023) align features of image-text pairs, DINOv2 (Oquab et al., 2024) employs self-supervised learning, and Pang et al. (2024) even uses Large Language Model (LLM) as a visual encoder. These models have shown strong capabilities in diverse visual tasks, including low-shot classification, open-vocabulary recognition, semantic segmentation, and visual question answering (VQA).
@@
Previous studies have reported similar conclusions as our studied representation biases when evaluating VFMs on different tasks (Ranzinger et al., 2024; Zong et al., 2024; Kar et al., 2024; Tong et al., 2024a; b). Therefore, several existing works have sought to leverage multiple VFMs to enhance downstream performance. Typically, they extract image features from multiple VFMs and then concatenate or fuse these features (Lin et al., 2023; Kar et al., 2024; Zong et al., 2024; Tong et al., 2024a; b; Man et al., 2024; Shlapentokh-Rothman et al., 2024). While this approach can improve visual encoding capability, it comes with a major drawback: running inference across multiple vision encoders drastically increases computational costs, as well as the memory and storage requirements due to large-scale parameters. Additionally, Zong et al. (2024) highlights that biased information from VFMs can lead to performance degradation when using a simple fusion method.
D.3 Knowledge Distillation of Vision Foundation Models
@@
Due to their large sizes and extensive training data requirements, vision foundation models are typically computationally intensive to train from scratch. Therefore, distilling VFMs into smaller models with knowledge distillation techniques has become a popular topic (Liu et al., 2024; Vemulapalli et al., 2024; Sun et al., 2023; Yang et al., 2024a; Zhang et al., 2024a). For example, several works (Zhang et al., 2023a; Zhou et al., 2023; Zhang et al., 2024b; Wang et al., 2023; Songa et al., 2024) have derived lightweight versions of SAM. More recently, research has explored distilling multiple VFMs into a single student model (Ranzinger et al., 2024; Shang et al., 2024; Sariyildiz et al., 2024), which is closely related to our approach. Besides the differences already discussed in the main paper, we further provide a detailed comparison in Table 15. RADIO uses strong teachers with substantially larger scales and capacities, for instance, DINOv2-g has 1,100M parameters. Theia also uses a ViT-H teacher with 632M parameters, double the size of ViT-L. Moreover, RADIO is trained on a dataset over 1,000 times larger than ImageNet-1k, even with larger input resolutions and a greater number of total samples seen. Despite the relatively lower training cost, our proposed SAK demonstrates superior outcomes on downstream tasks, exceeding both baselines by notable margins.
@@
Model RADIO (Ranzinger et al., 2024) Theia (Shang et al., 2024) SAK (Ours)
@@
D.4 Multi-Task Learning
@@
In multi-task dense prediction, most works focus on designing architecture, which can be further divided into encoder-focused and decoder-focused methods (Vandenhende et al., 2021). Encoder-focused methods develop shared encoders to extract features for different tasks, employing techniques like feature fusion (Misra et al., 2016; Ruder et al., 2019; Gao et al., 2019), attention (Liu et al., 2019), branched networks (Brüggemann et al., 2020; Guo et al., 2020; Vandenhende et al., 2019; Lu et al., 2017; Raychaudhuri et al., 2022), and mixture-of-experts (Liang et al., 2022; Chen et al., 2023; Ye & Xu, 2023a). Decoder-focused methods, on the other hand, design complicated decoders to extract task-specific features by modeling task interactions (Xu et al., 2018; Zhang et al., 2019; Zhou et al., 2020; Zhang et al., 2018; Brüggemann et al., 2021; Vandenhende et al., 2020; Ye & Xu, 2024; Xu et al., 2023c; b; Ye & Xu, 2023b; Zhang et al., 2021; Xu et al., 2022; Sirejiding et al., 2023; Xu et al., 2023a; Sirejiding et al., 2024b; Xin et al., 2024b; Sirejiding et al., 2024a; Lin et al., 2024; Huang et al., 2024a; Zhang et al., 2023b; Wang et al., 2024b). Additional topics include task-conditional models (Maninis et al., 2019; Kanakis et al., 2020; Sun et al., 2021; Lu et al., 2024c; b), efficient adaptation (Liu et al., 2022; Xin et al., 2024a; Agiza et al., 2024; Wang et al., 2024c), efficient computation (Neseem et al., 2023; Shoouri et al., 2023; Aich et al., 2023), regularization (Yang et al., 2023; Li et al., 2024), and generative modeling (Bao et al., 2022; Qiu et al., 2024).
Appendix E Additional Qualitative Results

### v2

@@ E.2 Vision Foundation Models
Vision Foundation Models (VFMs) are large-scale, general-purpose vision models trained on massive datasets, demonstrating exceptional performance across a wide range of downstream tasks, particularly when generalizing to unseen tasks and domains. Some VFMs are trained with specific tasks, such as ViT (Dosovitskiy et al., 2021), DeiT (Touvron et al., 2021), and Swin Transformer (Liu et al., 2021c) for image classification, Segment Anything Model (SAM) (Kirillov et al., 2023; Ravi et al., 2025) for promptable segmentation, and Depth Anything (Yang et al., 2024b; c) for monocular depth estimation. Others are trained with pretext tasks agnostic to downstream tasks. For instance, CLIP (Radford et al., 2021) and its derivatives (Zhai et al., 2023) align features of image-text pairs, DINOv2 (Oquab et al., 2024) employs self-supervised learning, and Pang et al. (2024) even uses Large Language Model (LLM) as a visual encoder. These models have shown strong capabilities in diverse visual tasks, including low-shot classification, open-vocabulary recognition, semantic segmentation, and visual question answering (VQA).
@@
Previous studies have reported similar conclusions as our studied representation biases when evaluating VFMs on different tasks (Ranzinger et al., 2024b; Zong et al., 2024; Kar et al., 2024; Tong et al., 2024a; b). Therefore, several existing works have sought to leverage multiple VFMs to enhance downstream performance. Typically, they extract image features from multiple VFMs and then concatenate or fuse these features (Lin et al., 2023; Kar et al., 2024; Zong et al., 2024; Tong et al., 2024a; b; Man et al., 2024; Shlapentokh-Rothman et al., 2024). While this approach can improve visual encoding capability, it comes with a major drawback: running inference across multiple vision encoders drastically increases computational costs, as well as the memory and storage requirements due to large-scale parameters. Additionally, Zong et al. (2024) highlights that biased information from VFMs can lead to performance degradation when using a simple fusion method.
E.3 Knowledge Distillation of Vision Foundation Models
@@
Due to their large sizes and extensive training data requirements, vision foundation models are typically computationally intensive to train from scratch. Therefore, distilling VFMs into smaller models with knowledge distillation techniques has become a popular topic (Liu et al., 2024c; Vemulapalli et al., 2024; Sun et al., 2023; Yang et al., 2024a; Zhang et al., 2024a). For example, several works (Zhang et al., 2023a; Zhou et al., 2023; Zhang et al., 2024b; Wang et al., 2023; Songa et al., 2024) have derived lightweight versions of SAM. More recently, research has explored distilling multiple VFMs into a single student model (Ranzinger et al., 2024b; Shang et al., 2024; Sariyildiz et al., 2024; Ranzinger et al., 2024a), which is closely related to our approach. Besides the differences already discussed in the main paper, we further provide a detailed comparison in Table 24. RADIO uses strong teachers with substantially larger scales and capacities, for instance, DINOv2-g has 1,100M parameters. Theia also uses a ViT-H teacher with 632M parameters, double the size of ViT-L. Moreover, RADIO is trained on a dataset over 1,000 times larger than ImageNet-1k, even with larger input resolutions and a greater number of total samples seen. Despite the relatively lower training cost, our proposed SAK demonstrates superior outcomes on downstream tasks, exceeding both baselines by notable margins.
@@
Model RADIO (Ranzinger et al., 2024b) Theia (Shang et al., 2024) SAK (Ours)
@@
E.4 Multi-Task Learning
@@
In multi-task dense prediction, most works focus on designing architecture, which can be further divided into encoder-focused and decoder-focused methods (Vandenhende et al., 2021). Encoder-focused methods develop shared encoders to extract features for different tasks, employing techniques like feature fusion (Misra et al., 2016; Ruder et al., 2019; Gao et al., 2019), attention (Liu et al., 2019), branched networks (Brüggemann et al., 2020; Guo et al., 2020; Vandenhende et al., 2019; Lu et al., 2017; Raychaudhuri et al., 2022), and mixture-of-experts (Liang et al., 2022; Chen et al., 2023; Ye & Xu, 2023a). Decoder-focused methods, on the other hand, design complicated decoders to extract task-specific features by modeling task interactions (Xu et al., 2018; Zhang et al., 2019; Zhou et al., 2020; Zhang et al., 2018; Brüggemann et al., 2021; Vandenhende et al., 2020; Ye & Xu, 2024; Xu et al., 2023c; b; Ye & Xu, 2023b; Zhang et al., 2021; Xu et al., 2022; Sirejiding et al., 2023; Xu et al., 2023a; Sirejiding et al., 2024b; Xin et al., 2024b; Sirejiding et al., 2024a; Lin et al., 2024; Huang et al., 2024a; Zhang et al., 2023b; Wang et al., 2024b). Additional topics include task-conditional models (Maninis et al., 2019; Kanakis et al., 2020; Sun et al., 2021; Lu et al., 2024c; b), efficient adaptation (Liu et al., 2022; Xin et al., 2024a; Agiza et al., 2024; Wang et al., 2024c), efficient computation (Neseem et al., 2023; Shoouri et al., 2023; Aich et al., 2023), regularization (Yang et al., 2023; Li et al., 2024a), and generative modeling (Bao et al., 2022; Qiu et al., 2024).
Appendix F Additional Qualitative Results